OpenStudy (anonymous):

Solve: x + 10 = −x + 12 8 −4 4 Question 7 (Multiple Choice Worth 3 points) Solve: 2(x + 6) − (x + 1) = 16 3 5 2 1 Question 8 (Multiple Choice Worth 3 points) Solve: 2 − x = 13 −11 11 −15 15 Question 9 (Multiple Choice Worth 3 points) Solve: 4(x + 1) = −12 −2 4 −3.25 −4

OpenStudy (nathan917):

1,5, idk for 8 but the last one is -2 @Blueskys might be able to help us with 8

OpenStudy (anonymous):

kk thanks man

OpenStudy (anonymous):

\[2-x=13\] \[2=x+13\] \[x=2-13\]\[x=-11\] hope it helps for question 8

OpenStudy (anonymous):

thanks @Blueskys

OpenStudy (nathan917):

Ty man.

OpenStudy (anonymous):

@R00nnYBraiNsbiG anytime

OpenStudy (anonymous):

can u help with 6 & 7

OpenStudy (nathan917):

i though i said it was 1 and 5 for 6 and 7 but to make sure i guess. im not a expert :D

OpenStudy (anonymous):

oh thats what u ment ok but @Blueskys how did u get -11

OpenStudy (anonymous):

2-x=13, we want x to be positive whenever possible because it looks prettier that way

OpenStudy (anonymous):

so we move it to the right hand side

OpenStudy (anonymous):

and theres no one 4 answer 6 so cnt b 1

OpenStudy (anonymous):

ill do 6 now

OpenStudy (anonymous):

kk thanks

OpenStudy (nathan917):

x +10= -x + 12 x+x+10= -x+x+12 (x cancels out) 2x+10=12 2x+10-10=12-10 (10's cancel) 2x=2 (divide by 2 on both sides) x=1 thats how i got 1 so..

OpenStudy (anonymous):

but 1 isnt a choice

OpenStudy (anonymous):

\[ x + 10 = −x + 12\]\[x+x=12-10\]\[2x=2\]\[x=1\] remember when moving from one side of the equal side to another, the number changes from positive to negative or the other way around

OpenStudy (anonymous):

theres only 8 -4 4 & 1/2

OpenStudy (anonymous):

might be misprinted

OpenStudy (anonymous):

so what do i choose 1/2 ?

OpenStudy (anonymous):

what was the question just to make sure

OpenStudy (anonymous):

Solve: x + 10 = −x + 12 8 −4 1/2 4

OpenStudy (anonymous):

same result \[x=1\]

OpenStudy (anonymous):

yea idk @radar please help

OpenStudy (nathan917):

Is there anymore to the question?

OpenStudy (anonymous):

my bad if I made a mistake

OpenStudy (anonymous):

no but in that huge blank it should b 1/2

OpenStudy (nathan917):

Yeah it might be 1/2 ill check it out

OpenStudy (radar):

If you have copied it correctly including the signs, Blueskys correctly solved it. x=1

OpenStudy (anonymous):

i did and thats wht i got is 1 but its not there

OpenStudy (nathan917):

*i said it was 1 first* then he did :D I'm never mentioned....

OpenStudy (nathan917):

It might be -4

OpenStudy (anonymous):

how did u get -4

OpenStudy (nathan917):

its -4 or 4 im not sure. ATM trying to solve it..

OpenStudy (radar):

The remainder do have a correct solution, but the first one is controversial.

OpenStudy (nathan917):

Simplifying x + 10 = -1x + 12 Reorder the terms: 10 + x = -1x + 12 Reorder the terms: 10 + x = 12 + -1x Solving 10 + x = 12 + -1x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add 'x' to each side of the equation. 10 + x + x = 12 + -1x + x Combine like terms: x + x = 2x 10 + 2x = 12 + -1x + x Combine like terms: -1x + x = 0 10 + 2x = 12 + 0 10 + 2x = 12 Add '-10' to each side of the equation. 10 + -10 + 2x = 12 + -10 Combine like terms: 10 + -10 = 0 0 + 2x = 12 + -10 2x = 12 + -10 Combine like terms: 12 + -10 = 2 2x = 2 Divide each side by '2'. x = 1 Simplifying x = 1 This question is messed up.... im just going to say its 1/2 best one that matches it..

OpenStudy (anonymous):

A PEOPLE I COPYIED IT WRONG ITS -1/2X+10=-X+12

OpenStudy (nathan917):

k

OpenStudy (anonymous):

@radar @Blueskys @nathan917

OpenStudy (radar):

close doesn't count in math, only in horseshoes and grenades!

OpenStudy (nathan917):

....

OpenStudy (radar):

Do you now get an answer that is closer like 4!

OpenStudy (anonymous):

YYEA

OpenStudy (nathan917):

i knew it was -4 or 4..

OpenStudy (radar):

Good luck and watch em close.

OpenStudy (nathan917):

Was Trying different equations.

OpenStudy (anonymous):

KK THANKS

OpenStudy (nathan917):

@radar @Blueskys Gj guys